OpenStudy (anonymous):

Solve the equation 4/5 y + 2 = 6 for y. Please show your work and solution.

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 } y + 2 = 6\]

OpenStudy (anonymous):

4/5y+2=6 4/5y=4 -2 both sides 4y=20 times 5 boths sides y=5 divide both sides by 4

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

THanks for medal